OpenStudy (okagurao):

What is the value of this expression when a=6 b=4 and c=8? 2a (pretend there is a line under 2a) 3b-c I will fan whoever gets the question right!

OpenStudy (okagurao):

Plllllzzz HAAAAAAAAAAAAAALLLLLLLLLLPPPP

OpenStudy (ccswims):

The answer is 3

OpenStudy (okagurao):

thats one of the answer choices once i am done with it (it wont be that long) I will fan you if its right

OpenStudy (ccswims):

because 2 x 6 = 12 then 3 x 4 = 12 - 8 = 4 and 12/4 = 3

OpenStudy (okagurao):

Thank you you deserve the title of LIFESAVER